for each given axis of symmetry and point on a parabola, determine another point on the parabola. a. axis of symmetry x=2; given point (0,5):
b. axis of symmetry x=-5; given point (-7,3):
c. axis of symmetry x=1/2; given point (-2,7):
d. axis of symmetry x=5; given point (10,0):

Answer by stanbon(75887)

for each given axis of symmetry and point on a parabola, determine another point on the parabola.
a. axis of symmetry x=2; given point (0,5)::::
(0,5) is 2 to the left of x=2; (4,5) is 2 to the right x =2
---------------------------------------------------------
b. axis of symmetry x=-5; given point (-7,3):
(-7,3) is 2 to the left of x=-5; (-3,3) is 2 to the right of x = -5
----------------------------------------------------------
c. axis of symmetry x=1/2; given point (-2,7):
(-2,7) is 2 1/2 to the left of x = 1/2; (3,7) is 2 1/2 to the right of x = 1/2
--------------------------------------------------------
d. axis of symmetry x=5; given point (10,0):
(10,0) is 5 to the right of x = 5; (0,0) is 5 to the left of x = 5
==============================
Cheers,
Stan H.